OpenStudy (t-dawg02):

what is 301/251 in simplest form?

rebeccaxhawaii (rebeccaxhawaii):

convert to a mixed number

OpenStudy (mathmale):

do 301 and 251 have any integer common factor? If so, find it, and use that factor to reduce the given fraction 301/251.

OpenStudy (t-dawg02):

i was thinking it was 50/251 but idk

rebeccaxhawaii (rebeccaxhawaii):

dont forget the one \[1\frac{ 50 }{ 251 }\neq \frac{ 50 }{ 251 }\]

rebeccaxhawaii (rebeccaxhawaii):

\[1\frac{ 50 }{ 251 }=\frac{ 301 }{ 251 }\]

OpenStudy (t-dawg02):

ok thx

rebeccaxhawaii (rebeccaxhawaii):

because